0 0
Add a comment Improve this question Transcribed image text
Answer #1

q = charge at the center of each sheet = 4 nC/3 = 1.33 x 10-9 C

for sheet 1 on left :

\theta = angle the line joining P to A makes with x-axis = tan-1(0.02/0.05) = 21.8

r = distance P to A = sqrt(0.052 + 0.022) = 0.054 m

Electric field is given as

E = k q/r2 = (9 x 109) (1.33 x 10-9)/(0.054)2 = 4.1 x 103 N/C

b)

By piece 2 :

r' = 0.02 m

Electric field is given as

E' = k q/r'2 = (9 x 109) (1.33 x 10-9)/(0.02)2 = 3 x 104 N/C

c)

By piece 3 :

\theta = angle the line joining R to A makes with x-axis = tan-1(0.02/0.05) = 21.8

r = distance P to A = sqrt(0.052 + 0.022) = 0.054 m

Electric field is given as

E = k q/r2 = (9 x 109) (1.33 x 10-9)/(0.054)2 = 4.1 x 103 N/C

d)

net electric field is given as

Enet = E Sin\theta + E' + E Sin\theta

Enet = 2E Sin\theta + E'

Enet = 2 (4.1 x 103) Sin21.8 + (3 x 104)

Enet = 3.3 x 104
Answer #1

q = charge at the center of each sheet = 4 nC/3 = 1.33 x 10-9 C

for sheet 1 on left :

\theta = angle the line joining P to A makes with x-axis = tan-1(0.02/0.05) = 21.8

r = distance P to A = sqrt(0.052 + 0.022) = 0.054 m

Electric field is given as

E = k q/r2 = (9 x 109) (1.33 x 10-9)/(0.054)2 = 4.1 x 103 N/C

b)

By piece 2 :

r' = 0.02 m

Electric field is given as

E' = k q/r'2 = (9 x 109) (1.33 x 10-9)/(0.02)2 = 3 x 104 N/C

c)

By piece 3 :

\theta = angle the line joining R to A makes with x-axis = tan-1(0.02/0.05) = 21.8

r = distance P to A = sqrt(0.052 + 0.022) = 0.054 m

Electric field is given as

E = k q/r2 = (9 x 109) (1.33 x 10-9)/(0.054)2 = 4.1 x 103 N/C

d)

net electric field is given as

Enet = E Sin\theta + E' + E Sin\theta

Enet = 2E Sin\theta + E'

Enet = 2 (4.1 x 103) Sin21.8 + (3 x 104)

Enet = 3.3 x 104